Vertical Analysis

Income statement information for Einsworth Corporation follows:

Sales $237,000
Cost of goods sold 78,210
Gross profit 158,790

Prepare a vertical analysis of the income statement for Einsworth Corporation. If required, round percentage answers to the nearest whole number.

Einsworth Corporation
Vertical Analysis of the Income Statement
Amount Percentage
Sales $237,000 %
Cost of goods sold 78,210 %
Gross profit $158,790 %
Answer #1
Einsworth Corporation
Vertical Analysis of the Income Statement
Amount Percentage
Sales $237,000 100%
Cost of goods sold 78,210 33%
Gross profit $158,790 67%

Cost of goods sold = 78210/237000 = 33%

Gross profit = 158,790 / 237000 = 67%